when my father was 35 i was 6.now he is twice as old as i am.how old am i??
Answer by sudheerb(5)

Before 'n' years from present, father age is 35 and son age is 6.
Present age of father is 35 + n.
Present age of son is 6 + n.
Present age of father is twice the present age of son.
(35 + n) = 2(6 + n)
35 + n = 2 x 6 + 2 x n
35 + n = 12 + 2n [subtract n from both sides of equation]
35 + n - n = 12 + 2n - n
35 = 12 + n. [subtract 12 from both sides of equation]
35 - 12 = 12 + n - 12
23 = n
n = 23.
Therefore present age of son is 6 + n = 6 + 23 = 29 years.